Operation Summary for November 12, 2010 - November 15, 2010 |
The Integrated Luminosity for the period from 11/8/10 to 11/15/10 was 49.02 pb-1 (inverse picobarns). NuMI reported receiving 7.18E18 protons on target during this same period. Friday Nov 12 The day shift began with the Tevatron (TeV) off for a D-Zero (D0) access, with the Collider Detector at Fermilab (CDF) not able to take data due to a detector problem, with Operations monitoring the Recycler stash (8.5E10), with the Antiproton Source (Pbar) stacking (70.3E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-992 experiments taking beam. During the day shift, D0 came out of access at 8:20 AM. Operations put the TeV into HEP shot setup and established store 8260 at 11:19 AM with an average initial luminosity (AIL) of 348.1E30. During the evening shift, Operations terminated store 8260 at 9:06 PM, put the TeV into shot setup, and established store 8261 at 10:24 PM with an AIL of 308.8E30. Saturday Nov 13 The midnight shift began with Operations monitoring store 8261 (209.03E30) and the stash (321.1E10), with Pbar stacking (17.6E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8261 at 6:02 AM. Operators prepared MI, the Pelletron, TeV, and the CDF and D0 collision halls for access at 6:12 AM. Operators had Booster prepared for access by 7:40 AM. During the day shift, most of the Accelerator Complex was in access. During the evening shift, Operations began sending beam to Pbar at 4:03 PM and to NuMI at 5:54 PM, but NuMI went down again at 7:28 PM due to a stuck LCW valve. Mechanical technicians repaired the valve and Operations had beam back to NuMI by 9:28 PM. Sunday Nov 14 The midnight shift began with the TeV off for CDF and D0 accesses, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Pelletron personnel completed their work and came out of access. During the day shift, CDF personnel completed their work and operators had their collision hall searched and secured by 1:42 PM. Operations reported at 3:03 PM that the TeV quenched during turn-on. Operators recovered the TeV and began its turn-on and checkout. During the evening shift, Operations put the TeV into shot setup at 10:10 PM and established store 8268 at 11:12 PM with an AIL of 318.7E30. Monday Nov 15 The midnight shift began with Operations monitoring store 8268 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations monitored the Accelerator Complex. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 11/15/2010 email Fermilab |
| Security, Privacy, Legal | |